You have collected data on several buildings. For each building, you are given the angle of the line of sight up to the top of the building, and the distance to the building. Calculate the height of each building.



15. Building 1 Angle 71o Distance 20 meters

A 89.26 m
B74.23 m
C25.62 m
D19.23 m
E 67.23 m
F 58.08 m

Answer by rothauserc(4718)

let h be the height of the building, then
tan(71) = h/20
h = tan(71)*20 = 58.08 m
F is the correct answer